OpenStudy (anonymous):

factor the expression. problem in comments. best answer i will become a fan.

OpenStudy (anonymous):

\[x^{2}+x-20\]

OpenStudy (anonymous):

Two numbers that multiply to give -20 and add to give 1 a*b=-20 a+b=1 a=5, b=-4 x^2+x-20=(x+a)(x+b)=(x+5)(x-4)

OpenStudy (anonymous):

im already a fan lol but thanks :)